# Support Graph Preconditioners for Off-Lattice Cell-Based Models

### v1

@@ 5 Numerical Benchmarks
In the previous sections, we extended support graph theory to block-structured matrices, introduced the collision graph, and argued for constructing preconditioners directly from the collision graph. Although we derived theoretical bounds on the eigenvalues of the preconditioned system, past experience with using the conjugate gradient method in finite precision arithmetic demonstrates that theoretical estimates often do not accurately predict performance in practice [26]. For this reason, we conduct benchmarking experiments to compare five preconditioning strategies: (1) no preconditioner, (2) the Jacobi preconditioner (i.e., the diagonal of $\Gamma$), (3) the block Jacobi preconditioner (i.e., the block diagonals of $\Gamma$), (4) the support graph MST preconditioner (as introduced in Section 3), and (5) a modified support graph MST preconditioner which uses the same block diagonals as the original matrix $\Gamma$.
5.1 Experimental Setup
@@
To test the preconditioning strategies, we generate two types of simulation-like test cases. Our experiments simulate of elastic spherical cells of radius $r=0.5$ (units are in 10s of $\mu m$). The Hertz contact model is used to resolve cell-cell contact areas and forces. We assume that the cell-substrate friction matrix is diagonal (i.e., $\Gamma^{cs}=\gamma_{\mathrm{med}}I$), and the cell-cell friction matrix is given by equation (2.1). We consider both cases in which $\gamma_{\parallel}=\gamma_{\perp}$ and $\gamma_{\parallel}\neq\gamma_{\perp}$. We find that cases in which $\gamma_{\parallel}\neq\gamma_{\perp}$ require more iterations for convergence. Typically the cell-cell friction coefficients are between 1–3 orders of magnitude greater than the cell-substrate friction coefficients [Galle2005]. We consider two types of spatial configurations for the cells:
  1. 1.
    Random placement: Cells are randomly placed in a 3D-sphere to simulate noisy cell configurations.
  2. 2.
    Hexagonal close packing: Each cell is surrounded by 12 other cells. Additionally, we apply normally distributed noise with mean zero and fixed standard deviation to vary the edge-vertex ratio.
The random configurations are created by randomly sampling locations in a 3D-sphere of given radius and only accepting locations that are a minimum distance from previously placed cells.
5.2 Benchmarking Results
In Fig. 3, we report the observed mean convergence behavior of the conjugate gradient method for the friction matrix of $N=10000$ randomly placed cells in a 3D-sphere. The relative error is computed by comparing the conjugate gradient solution to the solution obtained from a sparse direct method.
Rows:
The difference between the cell-substrate and cell-cell friction increases from single order of magnitude to three orders of magnitude across rows. We observe that more iterations are required for larger order of magnitude differences.
Columns:
The ratio of edges to vertices in the collision graph is increasing from left to right. We observe that the performance advantage of the support graph preconditioners becomes less pronounced as the edge-vertex ratio increases.
For low edge-vertex ratios, the support graph preconditioners result in approximately a $50\%$ iteration count reduction.
To explore the degradation in convergence rate of the support graph preconditioner as the edge-vertex ratio increases, we construct examples with high edge-vertex ratios. We generate cell configurations on three dimensional hexagonal lattices, so that each cell has 12 neighbors (equivalent to an edge-vertex ratio of six). The convergence results for these cases are reported in the last column of Fig. 4. In the first two rows of Fig. 4, there is a two order of magnitude difference between the cell-substrate friction and the cell-cell friction, while in the last row there is a three order of magnitude difference. We observe that the performance of the different preconditioners is comparable with the support graph preconditioner performing the worst. In the last row, for larger order of magnitude differences between cell-substrate and cell-cell friction, this difference is more pronounced.
To see whether we recover the support graph preconditioner performance as the edge-vertex ratio decreases, we add a normally distributed random variable with mean zero and standard deviation $\sigma$ to each cell location. As the standard deviation increases, the edge-vertex ratio decreases, and the support graph preconditioners perform best once more. In the cases with high edge-vertex ratios, the block Jacobi preconditioner outperforms the support graph preconditioner. This observation originally motivated us to consider the row-support graph preconditioners in which we use the block diagonals of the friction matrix together with the off-diagonals of the MST preconditioner.
Figure 4: Cells arranged on a three-dimensional hexagonal lattice with normal noise with mean zero and standard deviation $\sigma$ applied to each hexagonal cell position. In these examples we have $N=309$ cells. For each configuration we report the number of edges per vertex. As the edge to vertex ratio increases, the rate of convergence of both support graph preconditioners approaches that of the block-Jacobi preconditioners. For the hexagonal configuration without noise, where each cell has 12 neighbours the rate of convergence for all preconditioners is comparable. Here $\gamma_{\mathrm{med}}=3\times 10^{4}$.
6 Discussion
We have proposed efficient preconditioners for linear systems arising from off-lattice agent-based models. Using the notions of matrix-weighted graphs and block Laplacians, we extended support graph theory to this problem. By using maximum spanning trees, we obtained preconditioners that are computationally efficient to compute and factor, while significantly decreasing the condition number and iteration count in the conjugate gradient method. We proved bounds on the condition number that, while overly pessimistic in practice, demonstrate asymptotic stability. Additionally, we proved a bound for augmented spanning trees, but did not implement them because augmented trees lead to fill-in during factorization. In such cases, it may be beneficial to use the augmented MST preconditioner with an incomplete $LDL^{T}$ decomposition that completely ignores fill, however further research would be needed to develop a fill-reducing ordering. Finally, while we have not proved any bounds that theoretically justify including the entire block diagonal of the original matrix in the preconditioner, we provided numerical evidence of its potential helpfulness.
We benchmarked our proposed preconditioners and compared them to the identity, Jacobi, and block Jacobi preconditioners. Support graph preconditioners showed significant advantages in scenarios with low edge-vertex ratios, but their effectiveness decreased as this ratio increased. We observed that in situations where the support graph did not perform well, the block Jacobi preconditioner performed well. These findings led to the development of row-support graph preconditioners, combining block diagonal and support graph approaches. This combination resulted in a preconditioner which shares properties with both the block Jacobi and MST preconditioner. Our benchmarking supports this observation, as the row-support graph preconditioner performs well across various scenarios. Finally, our results highlight the importance of benchmarking, as theoretical bounds frequently do not predict practical performance in finite precision arithmetic.
What are the key takeaways for real-world implementations? Since the collision graph is a byproduct of the collision detection algorithm, it is logical to utilize it to define the system’s friction matrix. The collision graph allows for elegant implementations of the required linear algebra operations and the construction of an MST preconditioner. While our theoretical results give us asymptotic bounds on the eigenvalues of the preconditioned system, and numerical exploration indeed suggests a significant reduction in condition number, the observed number of required iterations in finite precision arithmetic paints a more complicated picture. The choice of preconditioner is problem-specific. The key considerations are the system’s edge-vertex ratio and magnitude difference between cell-substrate and cell-cell friction. The larger the difference between cell-substrate and cell-cell friction, the more useful preconditioning becomes. Similarly, the lower the edge-vertex ratio, the larger the reduction in required iterations when using the support graph preconditioner. Our results suggest that in practice, the edge-vertex ratio could be used to switch between support graph and block Jacobi preconditioning, or we could use the row-support graph preconditioner.
@@ Appendix A Positive Definite Off-Diagonal Blocks
There are two established ways of handling positive off-diagonal entries in the non-block case. The first is a reduction due to Gremban [17, Lemma 7.3] which is also described concisely in [35, Appendix A] and [28, Appendix A.2]. The validity of this reduction immediately transfers to the block case. Let $A$ be a block matrix whose off-diagonal blocks are either zero or definite (positive or negative), and for every row $i$,
@@
where ||$\cdot$|| ⋅ | leaves positive semidefinite matrices unchanged and negates negative semidefinite ones. This matrix is positive semidefinite by the proof from Lemma 2.4. We may call this a generalized block Laplacian, possibly originating from a similarly defined generalized matrix-weighted graph. Then $A=D+A^{(+)}+A^{(-)}$ where $D$ contains the diagonal blocks, $A^{(+)}$ the positive definite off-diagonal blocks, and $A^{(-)}$ the negative definite ones. To solve the system $A\mathbf{x}=\mathbf{b}$, we construct a $2n\times 2n$ block Laplacian system
@@ Appendix B Assorted Graph Algorithms
B.1 Directly Solving Systems from Trees
@@ 1. Step 1:
    Forward substitution. Define $\mathbf{z}=DL^{T}\mathbf{x}$ and solve $L\mathbf{z}=\mathbf{b}$, and proceed as in algorithm 3. The cost is $n-1$ matrix-vector multiplications of $d\times d$ matrices.
@@ 3. Step 3:
    Backward substitution. Solve $L^{T}\mathbf{x}=\mathbf{y}$. The cost is $n-1$ matrix-vector multiplications of $d\times d$ matrices.
@@
B.2 Matrix-Vector Product of a graph

### v2

@@ 5. Numerical Benchmarks
In the previous sections, we introduced the collision graph and block Laplacians, and we extended support graph theory to block-structured matrices, giving a class of preconditioners that can be contructed directly from the collision graph. Although we derived theoretical bounds on the eigenvalues of the preconditioned system, past experience with using the conjugate gradient method in finite precision arithmetic demonstrates that theoretical estimates often do not accurately predict performance in practice [27]. For this reason, we conduct benchmarking experiments to compare six preconditioning strategies: (1) no preconditioner, (2) the block Jacobi preconditioner (i.e., the block diagonals of $\Gamma$), (3) the symmetric block Gauss-Seidel preconditioner, (4) the MST preconditioner from Section 3.1, (5) the augmented MST preconditioner from Section 3.1, and (6) the block incomplete Cholesky preconditioner with no edge fill-in together with a diagonal shift $\alpha I$.
Our implementation begins with collision detection using an axis-aligned bounding box approach [39] that achieves $\mathcal{O}\left(n\log n\right)$ complexity with an event-based algorithm. This produces an active set of edges representing potential contacts, from which we compute contact forces, contact areas, and friction blocks according to equation (1) and the Hertz contact model. Our implementation uses a sweep and prune (SAP) algorithm without hierarchical subdivision, which limits our benchmarks to $n\leq 10^{5}$ cells where SAP remains efficient.
The resulting contact graph is converted to a block-structured BCSR (Block Compressed Sparse Row) format for efficient matrix-vector products, with reverse Cuthill-McKee (RCM) ordering to improve cache locality. The MST-preconditioner is solved using an $LDL^{T}$ decomposition directly on the graph structure. The augmented MST uses nested conjugate gradient iterations following the flexible CG framework [32]. Since we observed no iteration count improvements with augmentation, we did not pursue more efficient solution methods for this variant. All preconditioner construction operations (graph building, tree computation, and factorization) are cheaper than a single matrix-vector product, with friction blocks assembled during force calculation at no additional cost.
The solver is implemented in C++ using an expression-based linear algebra system with custom block-structured vectors that leverage BLAS routines. The preconditioner construction process, comprising Prim’s algorithm and tree assembly, is highly efficient, requiring less wall-clock time than a single matrix-vector product. After construction and factorization, solving with the MST preconditioner consumes 40–80% of the time needed for one matrix-vector product. Notably, approximately one-third of this solving time is spent permuting matrix entries between two orderings: the RCM ordering used for matrix-vector operations and the ordering from Prim’s algorithm used for factorization. This overhead suggests room for improvement through more sophisticated ordering strategies that simultaneously optimize both matrix-vector products and direct solves.
For our convergence benchmarks, we employ a rigorous testing framework. We work with known solutions to compute true relative errors and use Gauss-Radau error estimators [31] to monitor conjugate gradient convergence. To ensure fair comparisons across all benchmarks, we use single-threaded execution throughout.
5.1. Experimental Setup
@@
To test the preconditioning strategies, we simulate elastic spherical cells of radius $r=0.5$ (units are in 10s of $\mu m$) using the Hertz contact model for cell-cell contact areas and forces. We assume that the cell-substrate friction matrix is diagonal (i.e., $\Gamma^{cs}=\gamma_{\mathrm{med}}I$), and the cell-cell friction matrix is given by equation (1) with $\gamma_{\parallel}=2\times 10^{6}$ and $\gamma_{\perp}=8\times 10^{7}$, representing typical values where cell-cell friction coefficients are 1–3 orders of magnitude greater than cell-substrate friction [12]. We evaluate preconditioner performance across four distinct scenarios designed to test different geometric configurations and conditioning challenges:
  1. (1)
    Regular hexagonal lattice (baseline): Cells arranged in hexagonal close packing where each cell contacts 12 neighbors. We add positional noise of mean zero and standard deviation $0.3r$ to create varying cell-cell contact areas. We test with $\gamma_{\mathrm{med}}=3\times 10^{4}$ (well-conditioned) and $\gamma_{\mathrm{med}}=3\times 10^{3}$ (higher condition number).
  2. (2)
    Random packing: Cells are randomly placed within a spheroidal domain using a rejection sampling algorithm that enforces a minimum distance constraint to prevent overlap. This process generates the irregular connectivity patterns characteristic of biological tissues, avoiding the artifacts of regular lattices. For all simulations, we use $\gamma_{\mathrm{med}}=3\times 10^{4}$.
  3. (3)
    Dumbbell configuration: Two densely packed spheroids connected by a thin bridge of randomly placed cells. This tests preconditioner performance on graphs with bottleneck structures. We use $\gamma_{\mathrm{med}}=3\times 10^{4}$.
For the convergence behavior study (Figure 3), we use n=50 000n=50\,000 cells in each configuration and track the relative error reduction over conjugate gradient iterations. For the scaling study (Figure 4), we use the regular lattice configuration with $\gamma_{\mathrm{med}}=3\times 10^{4}$ and vary the problem size from $n=10^{2}$ to $10^{5}$ cells, measuring both iteration counts and wall-clock time.
All experiments use a relative tolerance of $10^{-5}$ for the conjugate gradient stopping condition, with true error computed using a known solution vector. Each data point represents the average of 5 independent random realizations.
5.2. Benchmarking Results
Figure 3 presents the convergence behavior of different preconditioners across four challenging geometric configurations with n=50 000n=50\,000 cells.
The MST preconditioner demonstrates superior performance across all test cases. On hexagonal lattices (top row), iteration reduction improves from $4.9\times$ at moderate condition number to $9.2\times$ at high condition numbers, significantly outperforming block Jacobi ($1.2\times$) and block Gauss-Seidel ($2.5\times$) for both moderate and high condition number scenarios. Performance remains strong on irregular geometries: $3.8\times$ reduction for random spheroidal packing and $8.5\times$ for the sparse dumbbell configuration, where traditional preconditioners struggle.
The augmented variant shows negligible improvement, likely because standard spanning trees already achieve low stretch for these graphs. Block Gauss-Seidel maintains consistent $2.5\times$ reduction across geometries, while diagonally shifted IC(0) degrades from $2.0\times$ to $1.3\times$ on sparse configurations.
Figure 4 demonstrates computational scaling from $n=10^{2}$ to $10^{5}$ cells. At production scale ($n=10^{5}$), the MST preconditioner achieves between $3.4\times$ and $4.4\times$ wall-clock speedup over standard methods, with preconditioner construction and application each requiring less effort than a single matrix-vector product. This efficiency, combined with favorable iteration count scaling (right panel), ensures performance advantages increase with problem size. Alternative methods scale poorly: the augmented MST preconditioner incurs overhead from nested iteration, while incomplete Cholesky performs poorly and requires tuning of the diagonal shift parameter.
6. Discussion
We have proposed efficient preconditioners for linear systems arising from off-lattice cell-based models. Using the notions of matrix-weighted graphs and block Laplacians, we extended support graph theory to this problem. By using MSTs, we obtained preconditioners that are efficient to compute and factor, while significantly decreasing the condition number, iteration count, and wall-clock time in the conjugate gradient method. We proved bounds on the condition number for both standard and augmented MSTs that, while overly pessimistic in practice, demonstrate asymptotic stability. Although augmented MST preconditioners have proven effective for mesh Laplacians, our implementation using nested conjugate gradient iterations showed no iteration count improvement for cell-based models. We hypothesize that the local interaction graphs from spherical cell packings already achieve sufficiently low stretch without augmentation, though other cell geometries or packing configurations might benefit from this technique.
We comprehensively benchmarked our proposed preconditioners against identity, block Jacobi, block Gauss-Seidel, block IC(0), and direct sparse solvers across problem sizes from $n=10^{2}$ to $10^{5}$. The MST preconditioner achieves the best wall-clock times for larger systems relevant to production simulations while remaining competitive with direct solvers for small systems. The magnitude of performance improvement varies with problem characteristics, particularly condition number and graph topology, rather than degrading systematically with connectivity. Indeed, our most substantial improvements occurred on challenging geometries: high difference in parallel and perpendicular cell-cell friction coefficients and near-disconnected configurations. This robust performance across diverse scenarios establishes the MST preconditioner as consistently advantageous for off-lattice cell-based models.
For practical implementation, several key insights emerge. The collision graph, naturally produced by collision detection algorithms, serves dual purposes: defining the system’s friction matrix and enabling direct construction of the MST preconditioner. This graph representation allows elegant implementations of the required linear algebra operations without intermediate matrix assembly. Construction of the MST preconditioner, including graph traversal, tree computation, and factorization, requires less computational effort than a single matrix-vector product, with friction blocks assembled during force calculation at no additional cost. The preconditioner solve requires only 40–80% of the time for one matrix-vector product, ensuring that iteration reductions translate directly to wall-clock performance gains.
@@ 6. Discussion
Future work will exploit the hierarchical spatial data structures already required for collision detection in large simulations ($n>10^{5}$) where SAP becomes inefficient without spatial subdivision [39]. These structures (octrees, Z-order, or Hilbert curve grids) partition cells into clusters of 300–1000 elements to maintain optimal SAP performance at each leaf. We envision constructing a hierarchical preconditioner where each leaf node maintains its own MST preconditioner for the local interaction subgraph. The full preconditioner consists of these independent local preconditioners applied in parallel, simply ignoring any interactions between regions. While this drops some edges from the preconditioner, it enables perfect parallelization with no inter-processor communication. The key insight is that the spatial partitioning required to maintain SAP efficiency also creates natural subdomains for embarrassingly parallel preconditioning, addressing both collision detection and linear solving bottlenecks simultaneously through a unified hierarchical approach.
@@ Appendix A Positive Definite Off-Diagonal Blocks
There are two established ways of handling positive off-diagonal entries in the non-block case. The first is a reduction due to Gremban [17, Lemma 7.3] which is also described concisely in [38, Appendix A] and [29, Appendix A.2]. The validity of this reduction immediately transfers to the block case. Let $A$ be a block matrix whose off-diagonal blocks are either zero or definite (positive or negative), and for every row $i$,
@@
where $|\cdot|$ leaves positive semidefinite matrices unchanged and negates negative semidefinite ones. This matrix is positive semidefinite by the proof from Lemma 2.4. We may call this a generalized block Laplacian, possibly originating from a similarly defined generalized matrix-weighted graph. Then $A=D+A^{(+)}+A^{(-)}$ where $D$ contains the diagonal blocks, $A^{(+)}$ the positive definite off-diagonal blocks, and $A^{(-)}$ the negative definite ones. To solve the system $A\mathbf{x}=\mathbf{b}$, we construct a $2n\times 2n$ block Laplacian system
@@ Appendix B Assorted Graph Algorithms
B.1. Directly Solving Systems from Trees
@@ 1. Step 1:
    Forward substitution. Define $\mathbf{z}=DL^{T}\mathbf{x}$ and solve $L\mathbf{z}=\mathbf{b}$, and proceed as in algorithm 3. The cost is n1n-1 matrix-vector multiplications of $d\times d$ matrices.
@@ 3. Step 3:
    Backward substitution. Solve $L^{T}\mathbf{x}=\mathbf{y}$. The cost is n1n-1 matrix-vector multiplications of $d\times d$ matrices.
@@
B.2. Matrix-Vector Product of a graph